# Photonic-Electronic Integrated Circuits for High-Performance Computing and AI Accelerator

### v1

@@ VI Outlook and Conclusion
In this review, we have outlined recent advancements in photonic-electronic integrated circuits for computing and AI tasks, covering the progress made at various levels, including devices, circuits, architecture, and systems, as well as cross-layer software-hardware co-design strategies. However, more technical challenges need to be addressed in the future to enable the particle application of photonic computing. Here, we point to possible research directions.
@@ VI-1 Device/new material innovation
As discussed in this review, the characteristics of optical components directly influence the performance of photonic-electronic computing circuits. Recent research in emerging materials and novel devices provide new opportunities for enhancing the performance of PICs for computing from the device level. For instance, heterogeneous modulators, leveraging such as III-V MOSCAP and 2-D materials [91, 100], demonstrate high modulation efficiency, a gigahertz-level cut-off frequency, and sub-picojoule-per-bit power consumption, allows a compact PIC layout with high energy efficiency. Additionally, the dataflow management in PIC-based computing, encompassing both digital and analog systems, mainly relies on electrical memory due to the absence of corresponding components in the photonic domain. Previous studies have shown that non-volatile materials and devices possess the capability for data storage, including pre-trained weight parameters [34, 113]. Significantly, non-volatile photonic memory, in contrast to conventional electronic memory, not only fulfills the role of data storage but also functions as a computational unit for ”in-memory” computing. Nevertheless, due to the limited setting frequency of current non-volatile materials, such PCM, scenarios necessitating dynamic data reading ans writing (particularly for high-performance digital processing units) call for high-speed, reconfigurable materials for optical memory.
@@
Beyond the innovation in optical components for computing reviewed in Section III-B3, optimizing other aspects of PICs also presents significant interest. Potential research directions encompass, but are not limited to (1) On-chip light sources to eliminate the complex fiber packaging and alignment processes; (2) integrated frequency combs as WDM sources [113]; (3) customized nonlinear units for on-chip implementation of programmable activation functions; and (4) efficient ADC/DAC and photodetector, specifically designed for optical computing architectures [191].
VI-2 Scale to Large Models and Advanced tasks
@@
Photonic computing still faces challenges concerning scalability when moving to support large models. Several directions can be explored. First, continue to explore domain-specific PTCs by trading universality for higher scalability, as indicated by previous research efforts [157, 158, 81, 115, 159, 117, 148]. Furthermore, there exists a substantial potential to further improve computational density by engineering tailored, compact photonic devices include multi-operand modulators and metasurface-based devices. Moreover, exploiting unique characteristics of photonics, such as wavelength, time, or mode-division multiplexing, can enable the reuse of hardware for a higher degree of parallel operations. Last, having a deep understanding the workload of evolving machine learning models is crucial to design the suitable photonic circuit design. For instance, attention-based Transformer models introduce dynamic matrix multiplication, challenging previous PTC designs optimized for CNNs. Recent advancements propose dynamically-operated PTC design [114] to handle dynamic matrix multiplication efficiently, ensuring optimal performance across diverse AI applications.
VI-3 On-chip ONN training protocol
@@
VI-4 Advanced photonics and electronics integration
The PICs for computing require on-chip interfacing with diverse electronic components, including signal conversion circuits and control systems, leading to substantial data communication between the electronic and photonic segments. To address this interconnect challenge, advanced integration methods can be employed, aiming to reduce interconnect complexity and associated losses, thereby improving overall system performance. Recent progress has shown promising results in monolithic fabrication processes [192, 193]. However, further breakthroughs, such as integration with more advanced CMOS nodes and 3-D electro-optical integration, are anticipated to enhance the capabilities of neuromorphic photonic chips.
@@ VI-5 Cross-layer co-design and Electronic-Photonic Design Automation (EPDA)
Cross-layer efforts open up additional opportunities to optimize the performance of ONNs, as discussed in Section V. With the growing complexity of photonic-electronic hardware platforms, exploring EPDA becomes critical to enhance productivity and efficiency, such as exploring automatic circuit layout generation and fast photonic circuit simulation [194].
@@ VI-7 Optical AI Software Stack
Developing specialized compilers and instruction sets tailored for photonic computing architectures is necessary to enable the integration of novel optical computing hardware into the mainstream AI software stack, as envisioned in Ref [195]. While existing deep learning frameworks such as PyTorch and TensorFlow can still serve as the front end, the compiler component must be adapted to generate machine code optimized for the unique photonic accelerators. This ensures efficient utilization of the optical hardware. Moreover, we anticipate open-source efforts in this direction to facilitate the progress of the optical AI software stack.
In conclusion, photonic-electronic integrated circuits stand out for their exceptional advantages in computing—in terms of low latency, high bandwidth, and energy efficiency— and exhibit a high potential to overcome the insurmountable bottlenecks of electronic computing. As innovative ideas and efforts continue to enhance the scalability and applicability of PICs across various dimensions, photonics is poised to become a key emerging technology in the long-term future to meet the increasing societal demand for high-performance computing and hardware accelerators for AI applications.

### v2

@@ VI Outlook and Conclusion
In this review, we have outlined recent advancements in photonic-electronic integrated circuits for computing and AI tasks, covering the progress made at various levels, including the device, circuit, architecture, and system level, as well as progress in cross-layer software-hardware co-design strategies. However, more technical challenges need to be addressed in the future to enable the practical application of photonic computing. Here, we point to possible research directions.
@@ VI-1 Device/new material innovation
As highlighted in this review, the characteristics of optical components directly impact the performance of photonic-electronic computing systems. Recent research in emerging materials and novel devices provide new opportunities for enhancing the performance of PICs from the device level. For instance, heterogeneous modulators, leveraging technologies such as III-V MOSCAP and 2-D materials [98, 111], demonstrate high modulation efficiency, gigahertz-level cut-off frequencies, and sub-picojoule-per-bit power consumption, allowing a compact layout with high energy efficiency. Additionally, dataflow management in PIC-based computing, which includes both digital and analog systems, primarily depends on electrical memory due to the lack of equivalent components in the photonic domain. Previous studies have demonstrated that non-volatile materials and devices have the capability to store data, including pre-trained weight parameters [32, 124]. Significantly, non-volatile photonic memory, in contrast to conventional electronic memory, serves not only as a data storage solution but also functions as a computational unit for ‘in-memory’ computing. Nevertheless, due to the limited switching frequency of current non-volatile materials, such as PCMs, scenarios requiring dynamic data reading and writing (particularly in high-performance digital processing units) demand high-speed, reconfigurable materials for optical memory.
@@
Beyond the innovation in optical components for computing reviewed in Section III-B3, optimizing other aspects of PICs also presents significant interest. Potential research directions encompass, but are not limited to (1) on-chip light sources to eliminate the complex fiber packaging and alignment processes; (2) integrated frequency combs as WDM sources [124]; (3) customized nonlinear units for on-chip implementation of programmable activation functions; and (4) efficient ADC/DACs and photodetectors, specifically designed for optical computing architectures [196].
VI-2 Advanced photonics and electronics integration
PICs for computing require on-chip interfacing with diverse electronic components, including signal conversion circuits and control systems, leading to substantial data communication between the electronic and photonic parts. Currently, wirebonding is a widely adopted strategy for E-O interconnects in academia. This method is particularly convenient and cost-effective for small-scale networks with few active devices, as it requires only the routing of driving signals to the edges of photonic chips, followed by connections to peripheral electronic units via wire bonds. However, as PICs scale up, managing metal wire routing within constrained chip areas presents a challenge. The beachfront bottleneck of photonic chips and electrical control units restricts the number of connections that can be implemented due to the limited chip perimeter. Furthermore, wirebonding can constrain the bandwidth of high-frequency devices owing to impedance mismatch.
To address this interconnect challenge, advanced integration methods can be employed, aiming to reduce interconnect complexity and improve overall system performance [197]. Flip-chip bonding directly connects two dies by soldering or using conductive bumps to align matching electrical pads on their surfaces. This approach allows flexible floorplanning and pad placement beyond chip edges, enabling higher interconnect density and reducing parasitic impedance. Given the fabrication process and chip sizes, typical integration strategies involve flip-chip bonding single or multiple heterogeneous dies (such as ADC/DAC, memory, microcontroller, FPGA, ASIC, etc.) onto a photonics chip which also serves as an interposer. However, this approach will invariably introduce large temperature excursions on the PIC varying with the workload. To address this issue, one strategy involves developing temperature-resilient photonic platforms [198]. Additionally, the heat spreader and through-silicon via (TSV) technique enables more flexible thermal management within chips [199, 200, 33, 201].
Monolithic fabrication processes that integrate both photonic and electronic components on a single substrate also have shown promising results [202, 203]. While this technology enables higher integration levels, bandwidth, and energy efficiency, significant potential remains for the optimization of fabrication processes, improving yield, and the development of comprehensive PDKs. Besides, further breakthroughs, such as integration with more advanced CMOS nodes and 2.5-D/3-D electro-optical integration, are anticipated to enhance the performance of photonic chips.
VI-3 Scale to large models and advanced tasks
@@
Photonic computing still faces challenges concerning scalability when moving to support large models. Several directions can be explored. First, the continued exploration of domain-specific PTCs which trade universality for higher scalability, as indicated by previous research efforts [165, 89, 127, 166, 129, 158]. Second, there exists a substantial potential to further improve computational density by engineering tailored, compact photonic devices, such as multi-operand modulators and metasurface-based devices. Third, the exploitation of the unique characteristics of photonics, such as wavelength, time, or mode-division multiplexing, can enable the reuse of hardware for a higher degree of parallel operations. Lastly, the development of a deep understanding of the workload of evolving machine learning models is crucial for designing suitable PICs. For instance, attention-based Transformer models introduce dynamic matrix multiplication, challenging previous PTC designs optimized for CNNs. Recent advancements propose dynamically-operated PTC design to handle dynamic matrix multiplication efficiently, ensuring optimal performance across diverse AI applications.
VI-4 On-chip ONN training protocol
@@ VI-5 Cross-layer co-design and electronic-photonic design automation (EPDA)
Cross-layer efforts open up additional opportunities to optimize the performance of ONNs, as discussed in Section V. With the growing complexity of photonic-electronic hardware platforms, exploring EPDA becomes critical for enhancing productivity and efficiency, such as exploring automatic circuit layout generation and fast photonic circuit simulation [204].
@@ VI-7 Optical AI software stack
Developing specialized compilers and instruction sets tailored for photonic computing architectures is necessary to enable the integration of novel optical hardware into the mainstream AI software stack, as envisioned in Ref [205]. While existing deep learning frameworks such as PyTorch and TensorFlow can still serve as the front end, the compiler component must be adapted to generate machine code optimized for the unique photonic accelerators. This ensures efficient utilization of the optical hardware. Moreover, we anticipate open-source efforts in this direction to facilitate the progress of the optical AI software stack.
In conclusion, photonic-electronic integrated circuits stand out for their exceptional advantages in computing—in terms of low latency, high bandwidth, and energy efficiency—and exhibit a high potential to overcome the insurmountable bottlenecks of electronic computing. Nevertheless, whether for digital or analog computing, the PIC scale of reported work remains limited, and further improvements in scalability are essential. Continuous research should aim to enhance throughput via innovative devices, architectural improvements, specialized training algorithms, and hardware-software co-design strategies. Additionally, efforts should focus on reducing power consumption and the costs associated with the E-O interface, in order to rival the performance of state-of-the-art electronic computing systems. This ambition, however, does not imply that the objective of optical computing is to surpass digital-electronic processors in all metrics and replace them. From an industry perspective, optical computing needs to identify niches where it excels over its electronic counterparts. A promising short-term application involves leveraging the high parallelism of optical computing to develop MVM processing units and accelerators for neural networks, an area that has been extensively studied. If the PIC scale can be well aligned with its tasks (including considerations for multiplexing), and use efficient E-O interfaces, then operation at a rate of one MVM per nanosecond could be achieved and offer significant advantages in high-throughput applications. With sustained innovation and effort, integrated photonics is poised to become a pivotal emerging technology, satisfying the escalating societal demand for high-performance computing and hardware acceleration in AI applications over the long term.